# C-Pack: Packaged Resources To Advance General Chinese Embedding

### v2

@@
Abstract
We introduce C-Pack, a package of resources that significantly advance the field of general Chinese embeddings. C-Pack includes three critical resources. 1) C-MTEB is a comprehensive benchmark for Chinese text embeddings covering 6 tasks and 35 datasets. 2) C-MTP is a massive text embedding dataset curated from labeled and unlabeled Chinese corpora for training embedding models. 3) C-TEM is a family of embedding models covering multiple sizes. Our models outperform all prior Chinese text embeddings on C-MTEB by up to +10% upon the time of the release. We also integrate and optimize the entire suite of training methods for C-TEM. Along with our resources on general Chinese embedding, we release our data and models for English text embeddings. The English models achieve state-of-the-art performance on the MTEB benchmark; meanwhile, our released English data is 2 times larger than the Chinese data. All these resources are made publicly available at https://github.com/FlagOpen/FlagEmbedding.
1 Introduction
Text embedding is a long-standing topic in natural language processing and information retrieval. By representing texts with latent semantic vectors, text embedding can support various applications, e.g., web search, question answering, and retrieval-augmented language modeling Karpukhin et al. (2020); Lewis et al. (2020); Guu et al. (2020). The recent popularity of large language models (LLMs) has made text embeddings even more important. Due to the inherent limitations of LLMs, such as world knowledge and action space, external support via knowledge bases or tool use is necessary. Text embeddings are critical to connect LLMs with these external modules Borgeaud et al. (2022); Qin et al. (2023).
Figure 1: The C-Pack resources to support general Chinese embedding.
@@ 1 Introduction
$\bullet$ Data. The development of general-purpose text embeddings puts forward much higher demands on the training data in terms of scale, diversity, and quality. To achieve high discriminative power for the embeddings, it may take more than hundreds of millions of training instances Izacard et al. (2021); Ni et al. (2021b); Wang et al. (2022b), which is orders of magnitude greater than typical task-specific datasets, like MS MARCO Nguyen et al. (2016) and NLI Bowman et al. (2015); Williams et al. (2017). Besides scale, the training data needs to be collected from a wide range of sources so as to improve the generality across different tasks Izacard et al. (2021); Wang et al. (2022b). Finally, the augmentation of scale and diversity will probably introduce noise. Thus, the collected data must be properly cleaned before being utilized for the training of embeddings Wang et al. (2022b).
@@
$\bullet$ Training. The training of general-purpose text embeddings depends on two critical elements: a well-suited backbone encoder and an appropriate training recipe. While one can resort to generic pre-trained models like BERT Devlin et al. (2018) and T5 (Raffel et al., 2020), the quality of text embedding can be substantially improved by pre-training with large-scale unlabeled data Izacard et al. (2021); Wang et al. (2022b). Further, instead of relying on a single algorithm, it takes a compound recipe to train general-purpose text embedding. Particularly, it needs embedding-oriented pre-training to prepare the text encoder Gao and Callan (2021), contrastive learning with sophisticated negative sampling to improve the embedding’s discriminability Qu et al. (2020), and instruction-based fine-tuning Su et al. (2022); Asai et al. (2022) to integrate different representation capabilities of text embedding.
@@
$\bullet$ Benchmark. Another pre-requisite condition is the establishment of proper benchmarks, where all needed capabilities of text embeddings can be comprehensively evaluated. BEIR Thakur et al. (2021) provides a collection of 18 to evaluate the embedding’s general performances on different retrieval tasks, e.g., question answering and fact-checking. Later, MTEB Muennighoff et al. (2022a) proposes a more holistic evaluation of embeddings and extends BEIR. It integrates 56 datasets, where all important capabilities of text embeddings, like retrieval, ranking, clustering, etc., can be jointly evaluated.
@@
Altogether, the development of general-purpose text embedding needs to be made on top of a mixture of driving forces, from data, and encoder models, to training methods and benchmarking. In recent years, continual progress has been achieved in this field, such as work from Contriever Izacard et al. (2021), E5 Wang et al. (2022b), and OpenAI Text Embedding Neelakantan et al. (2022). However, most of these works are oriented to the English world. In contrast, there is a severe shortage of competitive models for general Chinese embedding due to a series of limitations: there are neither well-prepared training resources nor suitable benchmarks to evaluate the generality.
@@
$\bullet$ C-MTEB (Chinese Massive Text Embedding Benchmark). The benchmark is established as a Chinese extension of MTEB. [^1] https://huggingface.co/spaces/mteb/leaderboard C-MTEB collects 35 public-available datasets belonging to 6 types of tasks. Thanks to the scale and diversity of C-MTEB, all major capabilities of Chinese embeddings can be reliably measured, making it the most suitable benchmark to evaluate the generality of Chinese text embedding.
$\bullet$ C-MTP (Chinese Massive Text Pairs). We create a massive training dataset of 100M text pairs, which integrates both labeled data and unlabeled data curated from Wudao Yuan et al. (2021), one of the largest corpora for pre-training Chinese language models. C-MTP is not only large and diverse but also cleaned to ensure the data quality.
$\bullet$ C-TEM (Chinese Text Embedding Models). We provide a family of well-trained models for Chinese general text embeddings. There are three optional model sizes: small (24M), base (102M), and large (326M), which present users with the flexibility to trade off efficiency and effectiveness. Our models make a big leap forward in generality: C-TEM outperforms all previously Chinese text embedding models on all aspects of C-MTEB by large margins. Besides being directly applicable, C-TEM can also be fine-tuned with additional data for better task-specific performances.
@@
In summary, C-Pack provides a go-to option for people’s application of general-purpose Chinese text embedding. It substantially advances the training and evaluation, laying a solid foundation for the future development of this field.
@@
Figure 2: Overview of C-Pack. C-MTEB is a benchmark for Chinese text embeddings. C-MTP is a large-scale Chinese embedding training dataset. C-TEM are state-of-the-art Chinese embedding models. The training recipe is shown at the bottom.
2 C-Pack
@@
In this section, we first introduce the resources in C-Pack: the benchmark C-MTEB, the training data C-MTP, and the model class C-TEM. Then, we discuss the training recipe, which enables us to train the state-of-the-art models for general Chinese embedding based on the offered resources.
2.1 Benchmark: C-MTEB
@@
In particular, we collect a total of 35 public datasets, all of which can be used to evaluate Chinese text embeddings. The collected datasets are categorized based on the embedding’s capability they may evaluate. There are 6 groups of evaluation tasks: retrieval, re-ranking, STS (semantic textual similarity), classification, pair classification, and clustering, which cover the main interesting aspects of Chinese text embeddings. Note that there are multiple datasets for each category. The datasets of the same category are collected from different domains and complementary to each other, therefore ensuring the corresponding capability to be fully evaluated.
The nature of each task and its evaluation metric are briefly introduced as follows.
$\bullet$ Retrieval. The retrieval task is presented with the test queries and a large corpus. For each query, it finds the Top-$k$ similar documents within the corpus. The retrieval quality can be measured by ranking and recall metrics at different cut-offs. In this work, we use the setting from BEIR Thakur et al. (2021), using NDCG@10 as the main metric.
$\bullet$ Re-ranking. The re-ranking task is presented with test queries and their lists of candidate documents (one positive plus N negative documents). For each query, it re-ranks the candidate documents based on the embedding similarity. The MAP score is used as the main metric.
$\bullet$ STS (Semantic Textual Similarity). The STS (Agirre et al., 2012, 2013, 2014, 2015, 2016) task is to measure the correlation of two sentences based on their embedding similarity. Following the original setting in Sentence-BERT Reimers and Gurevych (2019), the Spearman’s correlation is computed with the given label, whose result is used as the main metric.
$\bullet$ Classification. The classification task re-uses the logistic regression classifier from MTEB Muennighoff et al. (2022a), where the provided label is predicted based on the input embedding. The average precision is used as the main metric.
$\bullet$ Clustering. The clustering task is to group sentences into meaningful clusters. Following the original setting in MTEB Muennighoff et al. (2022a), it uses the mini-batch k-means method for the evaluation, with batch size equal to 32 and k equal to the number of labels within the mini-batch. The V-measure score is used as the main metric.
Finally, the embedding’s capability on each task is measured by the average performance of all datasets for that task. The embedding’s overall generality is measured by the average performance of all datasets in C-MTEB.
dataset
C-MTP (unlabeled)
C-MTP      (labeled)
source
Wudao, CSL, XLSUM-Zh, Amazon-Review-Zh, CMRC, etc.
[…]
size
[…]
Table 1: Composition of C-MTP
2.2 Training Data: C-MTP
We curate the largest dataset C-MTP for the training of general Chinese embedding. The paired texts constitute the data foundation for the training of text embedding, e.g., a question and its answer, two paraphrase sentences, or two documents on the same topic. To ensure the generality of the text embedding, the paired texts need to be both large-scale and diversified. Therefore, C-MTP is collected from two sources: the curation of massive unlabeled data, a.k.a. C-MTP (unlabeled); and the comprehensive collection of labeled data, a.k.a. C-MTP (labeled). The data collection process is briefly introduced as follows.
@@ 2.2 Training Data: C-MTP
$\bullet$ C-MTP (unlabeled). We look for a wide variety of corpora, where we can extract rich-semantic paired structures from the plain text, e.g., paraphrases, title-body. Our primary source of data comes from open web corpora. The most representative one is the Wudao corpus Yuan et al. (2021), which is the largest well-formatted dataset for pre-training Chinese language models. For each of its articles, we extract (title, passage) to form a text pair. Following the same recipe, we also collect such text pairs from other similar web content like Zhihu, Baike, news websites, etc. Aside from the open web content, we also explore other public Chinese datasets to extract text pairs, such as CSL (scientific literature), Amazon-Review-Zh (reviews), Wiki Atomic Edits (paraphrases), CMRC (machine reading comprehension), XLSUM-Zh (summarization), etc. The paired structures are obvious in these datasets, which are directly extracted for the augmentation of C-MTP (unlabeled).
@@
The text pairs curated from the web and other public sources are not guaranteed to be closely related. Therefore, data quality can be a major concern. In our work, we use a simple strategy to filter the data before adding it to C-MTP (unlabeled). Particularly, we use a third-party model: Text2Vec-Chinese [^2] https://huggingface.co/GanymedeNil to score the strength of relation for each text pair. We empirically choose a threshold of 0.43, and drop the samples whose scores are below the threshold. With such an operation, there are 100 million text pairs filtered from the unlabeled corpora. Despite the simplicity, we find that it effectively removes the irrelevant text pairs when manually reviewing samples and leads to strong empirical performances for the models trained on C-MTP (unlabeled).
[…]

### v3

# C-Pack: Packed Resources For General Chinese Embeddings

@@
Abstract.
We introduce C-Pack, a package of resources that significantly advance the field of general text embeddings for Chinese. C-Pack includes three critical resources. 1) C-MTP is a massive training dataset for text embedding, which is based on the curation of vast unlabeled corpora and the integration of high-quality labeled corpora. 2) C-MTEB is a comprehensive benchmark for Chinese text embeddings covering 6 tasks and 35 datasets. 3) BGE is a family of embedding models covering multiple sizes. Our models outperform all prior Chinese text embeddings on C-MTEB by more than +10% upon the time of the release. We also integrate and optimize the entire suite of training methods for BGE. Along with our resources on general Chinese embedding, we release our data and models for English text embeddings. The English models also achieve state-of-the-art performance on the MTEB benchmark; meanwhile, our released English data is 2 times larger than the Chinese data. Both Chinese and English datasets are the largest public release of training data for text embeddings. All these resources are made publicly available at https://github.com/FlagOpen/FlagEmbedding.
Text Embeddings, Training Data, Benchmark, Pre-trained Models
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 14–18, 2024; Washington, DC, USA.booktitle: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’24), July 14–18, 2024, Washington, DC, USAisbn: 979-8-4007-0431-4/24/07doi: 10.1145/XXXXXX.XXXXXXccs: Information systems Retrieval models and ranking
1. Introduction
Text embedding is a long-standing topic in natural language processing and information retrieval. By representing texts with latent semantic vectors, text embedding can support various applications, e.g., web search, question answering, and retrieval-augmented language modeling (Karpukhin et al., 2020; Lewis et al., 2020; Guu et al., 2020; Xiong et al., 2020). The recent popularity of large language models (LLMs) has made text embeddings even more important. Due to the inherent limitations of LLMs, such as world knowledge and action space, external support via knowledge bases or tool use is necessary. Text embeddings are critical to connect LLMs with these external modules (Borgeaud et al., 2022; Qin et al., 2023).
Figure 1. C-Pack presents 4 critical resources to support general Chinese embedding: C-MTEB (comprehensive evaluation benchmark), C-MTP (massive training data), BGE (powerful pre-trained models), the entire-suite of training recipe.
@@ 1. Introduction
$\bullet$ Data. The development of general-purpose text embeddings puts forward much higher demands on the training data in terms of scale, diversity, and quality. To achieve high discriminative power for the embeddings, it may take more than hundreds of millions of training instances (Izacard et al., 2021; Ni et al., 2021b; Wang et al., 2022b), which is orders of magnitude greater than typical task-specific datasets, like MS MARCO (Nguyen et al., 2016) and NLI (Bowman et al., 2015; Williams et al., 2017). Besides scale, the training data needs to be collected from a wide range of sources so as to improve the generality across different tasks (Izacard et al., 2021; Wang et al., 2022b). Finally, the augmentation of scale and diversity will probably introduce noise. Thus, the collected data must be properly cleaned before being utilized for the training of embeddings (Wang et al., 2022b).
@@
$\bullet$ Training. The training of general-purpose text embeddings depends on two critical elements: a well-suited backbone encoder and an appropriate training recipe. While one can resort to generic pre-trained models like BERT (Devlin et al., 2018) and T5 (Raffel et al., 2020), the quality of text embedding can be substantially improved by pre-training with large-scale unlabeled data (Izacard et al., 2021; Wang et al., 2022b). Further, instead of relying on a single algorithm, it takes a compound recipe to train general-purpose text embedding. Particularly, it needs embedding-oriented pre-training to prepare the text encoder (Gao and Callan, 2021), contrastive learning with sophisticated negative sampling to improve the embedding’s discriminability (Qu et al., 2020), and instruction-based fine-tuning (Su et al., 2022; Asai et al., 2022) to integrate different representation capabilities of text embedding.
@@
$\bullet$ Benchmark. Another pre-requisite condition is the establishment of proper benchmarks, where all needed capabilities of text embeddings can be comprehensively evaluated. BEIR (Thakur et al., 2021) provides a collection of 18 to evaluate the embedding’s general performances on different retrieval tasks, e.g., question answering and fact-checking. Later, MTEB (Muennighoff et al., 2022a) proposes a more holistic evaluation of embeddings and extends BEIR. It integrates 56 datasets, where all important capabilities of text embeddings, like retrieval, ranking, clustering, etc., can be jointly evaluated.
@@
Altogether, the development of general-purpose text embedding needs to be made on top of a mixture of driving forces, from data, and encoder models, to training methods and benchmarking. In recent years, continual progresses have been achieved in this field, such as Contriever (Izacard et al., 2021), E5 (Wang et al., 2022b), GTR (Ni et al., 2021b), and OpenAI Text Embedding (Neelakantan et al., 2022). Nevertheless, most of these models are dedicated to the English-centric scenarios. In contrast, there is a shortage of competitive models for general Chinese embedding. What is worse, the development of general Chinese embedding is severely constrained in many aspects: there are neither well-prepared training resources nor suitable benchmarks to evaluate the generality. [^1] This situation has been substantially improved since our work. New methods are continually developed on top of the benchmark, data, and models from C-Pack.
@@
$\bullet$ C-MTEB (Chinese Massive Text Embedding Benchmark). The benchmark is established as a Chinese extension of MTEB. [^2] https://huggingface.co/spaces/mteb/leaderboard C-MTEB collects 35 public-available datasets belonging to 6 types of tasks. We set up the unified testing protocols so that different embeddings can be evaluated on fair ground. We also develop the evaluation pipeline which significantly makes ease for the evaluation process. Thanks to the scale and diversity of C-MTEB, all major capabilities of Chinese embeddings can be reliably measured, making it the most suitable benchmark to evaluate the generality of Chinese text embedding.
$\bullet$ C-MTP (Chinese Massive Text Pairs). We create a massive training dataset of 100M text pairs. The majority of our dataset is curated from the massive web corpora, such as Baike (Wikipedia-style webs in Chinese), Zhihu (a major Chinese social media), major News Websites in Chinese. We extract the semantically related text pairs leveraging the rich-structured information within the data, such as title-to-document, subtitle-to-passage, question-to-answer, question-to-similar-question, etc. The extracted data is further cleaned for the massive weakly supervised training of the text embeddings. We also integrate diverse labeled datasets, which presents high-quality supervision signals for the final refinement of text embeddings. Besides, considering that there is no public available dataset for general English text embeddings either, we curate another massive dataset for English with the same method, which consists of 200M text pairs.
$\bullet$ BGE (BAAI General Embeddings). We provide a family of well-trained models for Chinese general text embeddings. There are three optional model sizes: small (24M), base (102M), and large (326M), which present users with the flexibility to trade off efficiency and effectiveness. Our models make a big leap forward in generality: BGE outperforms all previously Chinese text embedding models on all aspects of C-MTEB by large margins. Besides being directly applicable, BGE can also be fine-tuned with additional data for better downstream performances. The releasing of these powerful models substantially contributes to critical applications, such as search, question answering, and retrieval-augmented generation.
@@
Our project enjoys a widespread popularity in technique communities, like HuggingFace and Github. Remarkably, according to the up-to-date statics (2024-04), BGE model series have received more than 20 million downloads from HuggingFace since its release on 2023-08, making it one of the most popular embedding models in the world. They have been integrated by the major RAG and text-embedding frameworks in the world, such as Langchain [^3] https://python.langchain.com/docs/integrations/text_embedding/bge_huggingface, LLamaIndex [^4] https://docs.llamaindex.ai/en/stable/examples/embeddings/huggingface.html, and Huggingface [^5] https://huggingface.co/docs/text-embeddings-inference/index. The code-base also receives nearly 5,000 stars on GitHub. So far, there have been over 100 submissions on C-MTEB, and it is widely recognized as the most popular and authoritative benchmark for Chinese text embeddings. It’s worth noting that our project is still fast growing, with new resources continually created and released to the public. In summary, C-Pack provides a go-to option for the development, evaluation, and application of general-purpose Chinese text embedding, which establishes a solid foundation for the advancement of this field.
The remaining part of this paper is organized as follows. We discuss the related works about general text embeddings in Section 2. We present a detailed introduction about C-Pack resources in Section 3. We make comprehensive empirical analysis for the value of C-Pack in Section 4. Finally, we conclude our work in Section 5.
@@
Figure 2. Overview of C-Pack. C-MTEB is a benchmark for Chinese text embeddings. C-MTP is a large-scale Chinese embedding training dataset. BGE are state-of-the-art Chinese embedding models. The training recipe is shown at the bottom.
2. Related Work
The importance of general text embedding is widely recognized, not only for its wide usage in typical applications, like web search and question answering (Karpukhin et al., 2020) but also due to its fundamental role in augmenting large language models (Lewis et al., 2020; Guu et al., 2020; Borgeaud et al., 2022; Izacard et al., 2022; Shi et al., 2023). Compared with the conventional task-specific methods, the general text embedding needs to be extensively applicable in different scenarios. In recent years, there has been a continual effort in this field, where a series of well-known works are proposed, like Contriever (Izacard et al., 2021), GTR (Ni et al., 2021b), sentence-T5 (Ni et al., 2021a), Sentence-Transformer (Reimers and Gurevych, 2019), E5 (Wang et al., 2022a), OpenAI text embedding (Neelakantan et al., 2022), etc. Although it remains an open problem, recent studies highlight the following important factors.
$\bullet$ Firstly, the training data is desired to be large-scale and diversified, from which the embedding model can learn to recognize different kinds of semantic relationships (Izacard et al., 2021; Ni et al., 2021b; Wang et al., 2022b; Neelakantan et al., 2022). It usually calls for the comprehensive and elaborate data curation from web corpora, such as online encyclopedia, QA platforms, news websites, and social media communities. Despite similar efforts made by the previous works, few of the curated datasets are made public available before C-Pack.
$\bullet$ Secondly, the embedding model must be scaled up in terms of both training and model size, as scaled-up text encoders are more generalizable across different application scenarios (Muennighoff, 2022; Ni et al., 2021b, a). Such an observation is in line with the conclusion for the importance of scaling LLMs (Hoffmann et al., 2022; Rae et al., 2021; Brown et al., 2020; Chowdhery et al., 2022; Srivastava et al., 2022; Gao et al., 2021a; Li et al., 2023a; Allal et al., 2023; Muennighoff et al., 2023b). Although large-scale training is affordable for many industrial organizations and companies, it remains a huge burden for the community users. The public release of BGE saves such a cost, benefiting both direct applications of text embeddings and further improvements for specific scenarios.
$\bullet$ Thirdly, the training recipe must be optimized through pre-training (Liu and Shao, 2022; Wang et al., 2022a), negative sampling (Izacard et al., 2021; Wang et al., 2022a), and multi-task fine-tuning (Su et al., 2022; Asai et al., 2022; Sanh et al., 2021; Wei et al., 2021; Muennighoff et al., 2022b, 2023a; Chung et al., 2022). In C-Pack, these operations are integrated, optimized, and pipelined, which significantly facilitates people’s reproduction and continual fine-tuning of BGE.
$\bullet$ Aside from the above factors, it is also critical to establish proper benchmarks to evaluate the generality of text embeddings. Unlike previous task-specific evaluations, like MSMARCO (Nguyen et al., 2016), SentEval (Conneau and Kiela, 2018), it is needed to substantially augment the benchmarks so as to evaluate the embedding’s performance for a wide variety of tasks. One representative work is made by BEIR (Thakur et al., 2021; Kamalloo et al., 2023), where the embeddings can be evaluated across different retrieval tasks. It is later extended by MTEB (Muennighoff et al., 2022a), where all major aspects of text embeddings can be comprehensively evaluated. However, no such works were done for the Chinese community before. By introducing C-MTEB, this limitation has been substantially conquered.
Given the above analysis, it can be concluded that the general text embedding is highly resource-dependent, which calls for a wide range of elements, such as datasets, models, and benchmarks. Thus, the creation and public release of the corresponding resources in C-Pack is crucially important.
3. C-Pack
@@
In this section, we first introduce the resources in C-Pack: the benchmark C-MTEB, the training data C-MTP, and the model class BGE. Then, we discuss the training recipe, which enables us to train the state-of-the-art models for general Chinese embedding based on the offered resources.
3.1. Benchmark: C-MTEB
C-MTEB is established for the comprehensive evaluation of the generality of Chinese embeddings (Figure 2). In the past few years, the community has put forward many datasets for text representation and language understanding tasks in Chinese, such as CMNLI (Xu et al., 2020), DuReader (He et al., 2017), T2Ranking (Xie et al., 2023). However, these datasets are independently curated, lacking a fair and shared ground to comprehensively evaluate the general capability of text embeddings. Therefore, we create C-MTEB, where the following important efforts are made: 1) the comprehensive collection of datasets which can be either directly utilized or repurposed for the evaluation of text embeddings, 2) the categorization of the datasets into different capability attributes of text embeddings, e.g., retrieval, similarity analysis, classification, etc., 3) the standardization of the evaluation protocols, 4) the establishment of evaluation pipelines.
In particular, we collect a total of 35 public datasets related to the evaluation of Chinese text embeddings (Briefed as Figure 1. Detailed specifications are presented in the Github Repository [^6] https://github.com/FlagOpen/FlagEmbedding/tree/master/C_MTEB). The collected datasets are categorized based on the embedding’s capability they may evaluate. There are 6 groups of evaluation tasks: retrieval, re-ranking, STS (semantic textual similarity), classification, pair classification, and clustering, which cover the main interesting aspects of Chinese text embeddings. Note that there are multiple datasets for each category. The datasets of the same category are collected from different domains and complementary to each other, therefore ensuring the corresponding capability to be fully evaluated.
The nature of each evaluation task and the evaluation metric are briefly introduced as follows.
$\bullet$ Retrieval. The retrieval task is presented with the test queries and a large corpus. For each query, it finds the Top-$k$ similar documents within the corpus. The retrieval quality can be measured by ranking and recall metrics at different cut-offs. In this work, we use the setting from BEIR (Thakur et al., 2021), using NDCG@10 as the main metric.
$\bullet$ Re-ranking. The re-ranking task is presented with test queries and their candidate documents (1 positive plus N negative documents). For each query, it re-ranks the documents based on the embedding similarity. The MAP score is used as the main metric.
$\bullet$ STS (Semantic Textual Similarity). The STS (Agirre et al., 2012, 2013, 2014, 2015, 2016) task is to measure the correlation of two sentences based on their embedding similarity. Following the original setting in Sentence-BERT (Reimers and Gurevych, 2019), the Spearman’s correlation is computed with the given label, whose result is used as the main metric.
$\bullet$ Classification. The classification task re-uses the logistic regression classifier from MTEB (Muennighoff et al., 2022a), where the average precision is used as the main metric.
$\bullet$ Clustering. The clustering task is to group sentences into meaningful clusters. Following the original setting in MTEB (Muennighoff et al., 2022a), it uses the mini-batch k-means method for the evaluation, with batch size equal to 32 and k equal to the number of labels within the mini-batch. The V-measure score is used as the main metric.
Figure 3. The evaluation pipeline of C-MTEB.
Finally, the embedding’s capability on each task is measured by the average performance of all datasets for that task. The embedding’s overall generality is measured by the average performance of all datasets in C-MTEB. We set up the standardized and compact pipeline for all tasks where different embedding models can be evaluated on a fair basis (Figure 3). FlagDRESModel is the wapper for the custmoized embedding model, which implements the encoding method for query and document. ChineseTaskList is the task list of C-MTEB. The evaluation result is written to the output folder, which can be directly submitted to the C-MTEB leaderboard [^7] https://huggingface.co/spaces/mteb/leaderboard.
3.2. Training Data: C-MTP
We curate the largest dataset C-MTP for the training of general Chinese embedding. The paired texts constitute the data foundation for the training of text embedding, e.g., a question and its answer, two paraphrase sentences, or two documents on the same topic. To ensure the generality of the text embedding, the paired texts need to be both large-scale and diversified. Therefore, C-MTP is collected from two sources. The majority of the data is based on the curation of massive unlabeled data, a.k.a. C-MTP (unlabeled), which presents 100 millions of paired texts. Meanwhile, a small portion is from the comprehensive integration of high-quality labeled data, a.k.a. C-MTP (labeled), which leads to about 1 million paired texts. The data collection process is briefly introduced as follows.
Table 1. Composition of C-MTP
dataset C-MTP (unlabeled) C-MTP      (labeled)
source Wudao, Zhihu, Baike, CSL, XLSUM-Zh, Amazon-Review-Zh, CMRC, etc. T2-Ranking, mMARCO-Zh, DuReader, NLI-Zh, etc.
@@ 3.2. Training Data: C-MTP
Figure 4. Creation of C-MTP.
@@
$\bullet$ C-MTP (unlabeled). We look for a wide variety of corpora, where we can extract rich-semantic paired structures from the plain text, e.g., paraphrases, title-body. Our primary source of data comes from open web corpora. The most representative one is the Wudao corpus (Yuan et al., 2021), which is the largest dataset of well-formatted articles for pre-training Chinese language models. For each its article, we extract structures, like (title, body), (sub-title, passage), to form a text pair. In addition, we collect data from other web content like Zhihu, Baike, news websites, which complement other forms of text pairs, especially (question, answer), (paraphrase titles), (paraphrase answers), etc. Aside from the open web content, we also explore other public Chinese data for more diverse text pairs, such as CSL (scientific literature), Amazon-Review-Zh (topic and its reviews), Wiki Atomic Edits (paraphrases), CMRC (machine reading comprehension), XLSUM-Zh (summarization), etc.
@@
The text pairs curated from the web and other public sources are not guaranteed to be closely related. Therefore, data quality can be a major concern. In our work, we make use of a compound data cleaning strategy to refine the raw data. Firstly, the whole data undergoes general filtering, which removes non-textual, duplicated, and malicous content. Secondly, the data is further processed by semantic filtering so as to ensure the text pairs are semantically related. In our work, we make use a third-party model: Text2Vec-Chinese [^8] https://huggingface.co/GanymedeNil to score the strength of relation for each text pair. We empirically choose a threshold of 0.43, and drop the samples whose scores are below the threshold. With such an operation, there are 100 million text pairs filtered from the unlabeled corpora. Despite the simplicity, we find that it effectively removes the irrelevant text pairs when manually reviewing samples and leads to strong empirical performances for the models trained on C-MTP (unlabeled).
$\bullet$ C-MTP (labeled). The labeled data is collected to further enhance to the training data. In our work, we integrate and re-purpose a diverse group of datasets, which covers different capabilities of the text embedding, like retrieval, ranking, similarity comparison, etc. Particularly, the following labeled datasets are included, T2-Ranking (Xie et al., 2023), DuReader (He et al., 2017; Qiu et al., 2022), mMARCO (Bonifacio et al., 2021), CMedQA-v2(Zhang et al., 2018), multi-cpr(Long et al., 2022), NLI-Zh [^9] https://huggingface.co/datasets/shibing624/nli_zh, cmnli(Xu et al., 2020) and ocnli(Xu et al., 2020). There are 838,465 paired texts in total, which contains diverse question-answering and paraphrasing patterns. Although it is much smaller than C-MTP (unlabeled), most of the data is curated from human annotation, thus ensuring a high credibility of relevance.